OpenStudy (anonymous):

Solve the system using elimination. 5x + 4y = 12 3x – 3y = 18

OpenStudy (gk99):

3x = 18 + 3y x = 6 + y so 5(6 + y) + 4y = 12 30 + 5y + 4y = 12 9y = -28 y= -18/9 y= -2

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (gk99):

np!